A large truck tire is inflated to a gauge pressure of 88 psi. The total area of one sidewall of the tire is 1,220 in2. What is the net outward force (in lb) on the sidewall because of the air pressure? (Enter the magnitude.)
1
Expert's answer
2019-04-11T09:52:12-0400

In accordance with the definition of a gauge pressure, the desired force is F = pS = 88 * 1220 = 107360 (lb).